OpenStudy (anonymous):

The three angles in a triangle always add up to 180°. If one angle in a triangle is 55° and the second is 4 times the third, what are the three angles? a. 55°, 100°, 25° b. 55°, 98°, 27° c. 55°, 101°, 24° d. 55°, 99°, 26° The three angles in a triangle always add up to 180°. If one angle in a triangle is 55° and the second is 4 times the third, what are the three angles? a. 55°, 100°, 25° b. 55°, 98°, 27° c. 55°, 101°, 24° d. 55°, 99°, 26° @Mathematics

OpenStudy (king):

a

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

let the third angle be x, 55+4x+x=180; 180-55=5x; 125=5x; x= 25; substitute x; 4*25 =100, so option a.